OpenStudy (anonymous):

let f(x)=x+5 and g(x) = x^2. Find (g f)(4)

OpenStudy (anonymous):

Possible answers: 9 81 20 1

OpenStudy (missinglink):

f(4) = 4 + 5 = 9; g(f(4)) = g(9) = 9^2 = 81. I think that is all there is to this problem, but the remember that (g f)(4) is not the same as (f g)(4).

OpenStudy (anonymous):

thanks!